On April 2, 2017, Victor, Inc. acquired a new piece of filtering equipment. The cost of the equipment was $400,000 with a residual value of $30,000 at the end of its estimated useful lifetime of 5 years. Assume that in its financial statements, Victor uses straight-line depreciation and the half-year convention. Depreciation recognized on this equipment in 2017 and 2018 will be:

Multiple Choice

$80,000 in 2017 and $64,000 in 2018.

$30,000 in 2017 and $74,000 in 2018.

$37,000 in 2017 and $74,000 in 2018.

$55,500 in 2017 and $64,000 in 2018.
Answer #1

Annual Depreciation expense = (Cost-salvage value)/Useful life = (400000-30000/5) = 74000

Under half year convention, it is assumed that assets is purchased in half year and in first year half year Depreciation would be recorded

So

2017 Depreciation expense = 74000*6/12 = 37000

2018 Depreciation expense = 74000

So answer is c) $37000 in 2017 and $74000 in 2018